# DPMambaIR: All-in-One Image Restoration via Degradation-Aware Prompt State Space Model

### v2

@@ III-D Degradation-aware Prompt State Space Model
To address the limitations of SSM-based image restoration—specifically the insufficient use of global context and the lack of effective degradation-aware mechanisms, we propose a novel Degradation-aware Prompt State Space Model (DP-SSM) to fully unlock the potential of SSM in All-in-One image restoration. Specifically, we propose a novel Degradation-aware Prompt Selective Scan, which incorporates degradation information into the State Space Model. This is achieved by leveraging a Degradation Embedding extracted from a pre-trained Degradation Extractor to provide supplementary degradation-related contextual information that the traditional SSM cannot capture. In addition, we modulate the parameter matrices in the state-space equations to establish a degradation-aware mechanism. This mechanism enables SSM to adapt to multi-task learning and handle task competition. To better recover local high-frequency details, we design a lightweight High-frequency Enhancement Block (HEB). This component ensures the preservation of high-frequency details and enhances the overall model performance.
@@
where $init()$ is an initialization method that does not depend on the input $x$. However, such parameterization does not incorporate degradation information, which motivates us to introduce a degradation-aware prompt mechanism to guide the state-space modeling.
@@
$\overline{B}_{dp}$ $=(\Delta_{dp}A)^{-1}(\exp(A)-I)\cdot\Delta_{dp}B_{dp},$
@@
Through this approach, the transformation matrix $A$ and input matrix $B$ gain degradation-aware capabilities, integrating degradation cues of various types and levels during historical information compression and input signal processing, making the model more task-friendly. The output matrix $C$ controls how historical information affects the output. In traditional SSM, the $i$-th pixel can only observe the i1i-1 preceding pixels, leading to a limited global receptive field. While existing methods address this with parallel multi-directional scanning, this increases computational costs and offers limited performance gains in low-level vision tasks [32]. Instead, we retain single-directional scanning and enhance global degradation cues via $C$, achieving efficient global awareness with lower computational overhead.
High-frequency Enhancement Block.
@@
All-in-One image restoration involves recovering image components distributed across different frequency bands. While our proposed Degradation-aware Prompt State Space Model enables dynamic adaptation to degradation characteristics and partially alleviates inter-task conflicts, it still exhibits a bias toward low-frequency restoration due to the inherent global modeling nature of Mamba. As shown in Fig. 5, All-in-One models tend to underperform in reconstructing high-frequency details such as edges and textures, compared to degradation-specific models. To address this limitation, we introduce an extremely lightweight High-Frequency Enhancement Block (HEB) as a complementary module.
@@ III-E Loss Function
where $O$ and $\hat{O}$ denote the model output and the ground truth, respectively. The parameters $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are balancing factors, which we set to 1, 0.5 and 0.001, respectively.
@@
TABLE I: Quantitative Comparison on Seven Datasets for All-in-One Image Restoration Task with State-of-the-Art Methods. The Best and Second-Best Results Are in Bold and Underline.
Method Lowlight Snow Rain Haze Jpeg Noise Blur Average PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM MPRNet 19.71 0.659 30.94 0.938 26.27 0.912 25.63 0.947 30.51 0.938 28.03 0.897 26.07 0.869 26.74 0.880 MIRNet 19.81 0.668 31.04 0.941 25.77 0.907 22.26 0.882 30.54 0.939 28.14 0.900 26.32 0.873 26.27 0.873 NAFNet 19.49 0.650 30.33 0.935 26.04 0.906 25.63 0.946 30.20 0.936 27.80 0.895 26.13 0.870 26.52 0.877 Restormer 19.99 0.670 31.67 0.944 26.91 0.919 26.18 0.953 30.56 0.939 28.17 0.890 26.62 0.880 27.16 0.886 MambaIR 19.81 0.664 30.43 0.935 25.52 0.901 25.90 0.944 30.38 0.937 27.91 0.895 26.07 0.869 26.57 0.878 PromptIR 19.90 0.669 31.52 0.943 26.80 0.918 26.13 0.945 30.57 0.939 28.18 0.901 26.36 0.874 27.07 0.884 IDR 19.68 0.663 31.16 0.941 26.39 0.914 26.11 0.948 30.51 0.939 28.10 0.899 26.29 0.873 26.89 0.883 NDR-Restore 19.79 0.635 30.00 0.902 26.00 0.813 27.00 0.943 30.19 0.872 27.78 0.786 25.77 0.791 26.64 0.820 OneRestore 19.77 0.657 29.95 0.931 25.37 0.900 27.74 0.962 30.10 0.934 27.65 0.890 25.90 0.866 26.64 0.877 MoCEIR 19.92 0.674 31.22 0.943 26.56 0.920 28.06 0.963 30.51 0.939 28.13 0.900 26.34 0.874 27.25 0.888 AdaIR 19.99 0.672 31.94 0.923 27.01 0.922 26.61 0.952 30.53 0.939 28.19 0.900 26.61 0.879 27.26 0.887 DPMambaIR 20.04 0.680 32.36 0.949 27.07 0.926 28.13 0.963 30.57 0.939 28.25 0.902 27.43 0.894 27.69 0.893 DPMambaIR-L 20.06 0.677 32.31 0.950 27.11 0.925 28.39 0.963 30.55 0.939 28.21 0.902 27.45 0.895 27.73 0.893
@@ IV-A Experimental Settings
Training Details. Building on previous works[39, 22, 32], we develop a Mamba-based baseline for image restoration. Our method is trained using the AdamW[41] optimizer with parameters $\beta_{1}=0.9$, $\beta_{2}=0.999$, and a weight decay of $1e^{-4}$ for 300,000 iterations. The initial learning rate of $3e^{-4}$ is gradually reduced to $1e^{-6}$ through cosine annealing. Following[39], we employ a progressive training strategy, cropping image patches of size $192\times 192$. Horizontal and vertical flips are applied for data augmentation. Our experiments are conducted on a single NVIDIA RTX A100 GPU and implemented using the PyTorch platform.
@@
Evaluation Metrics. Following previous works[39, 32], we use peak signal-to-noise ratio (PSNR) and structural similarity index measure (SSIM) to quantitatively compare different methods on the RGB channels.
@@ IV-B Quantitative Results
All-in-one Image Restoration. We conducted two versions of our method for comparison: the normal version (DPMambaIR) and a Large version (DPMambaIR-L). The normal version adopts the degradation extractor illustrated in Fig. 3(c), while the Large version uses the architecture shown in Fig. 3(b). For subsequent evaluations, we primarily adopt the normal version unless otherwise specified. We compare our method against eleven state-of-the-art image restoration approaches. As shown in Table I, for the All-in-One image restoration task, our method achieves the best PSNR and SSIM on the aforementioned mixed datasets, outperforming existing methods. Specifically, our method achieves a PSNR of 27.69dB and an SSIM of 0.893, surpassing AdaIR by 0.43dB and 0.011, respectively. Additionally, DPMambaIR-L achieves a PSNR of 27.73dB and an SSIM of 0.893, slightly outperforming the normal version.
@@ IV-C Qualitative Results
IV-D Ablation Study
We conduct ablation studies to evaluate our key designs: the degradation-aware prompts and High-frequency Enhancement Block (HEB) (Table IV), degradation extraction and utilization methods (Table V), and the global information supplementation strategy (Table VI).
Table IV shows the effectiveness of our core modules. The High-frequency Enhancement Block (HEB) alone improves the baseline by 0.29 dB PSNR by enhancing local details. Degradation-aware prompts ($P_{\Delta}$, $P_{B}$) bring a more significant gain of 0.64 dB, showing the benefit of adaptive modulation. Adding $P_{C}$ for global cues offers further slight improvement. The full model demonstrates the synergy between prompt-based modulation and frequency enhancement, achieving the best performance.
TABLE IV: Ablation Studies on Our Proposed Core Modules.
Method $P_{\Delta}$ $P_{B}$ $P_{C}$ HEB PSNR SSIM Baseline 26.96 0.884 (a) 27.25 0.887 (b) 27.60 0.892 (c) 27.67 0.893 (d) 27.69 0.893
@@ IV-D Ablation Study
Table V presents the results for degradation extraction and utilization methods. For degradation extraction, we compare three approaches: (1) “Classes”, which provides manually specified degradation types and can be regarded as a 100% accurate non-blind setting; (2) a pre-trained degradation classifier from OneRestore [9]; and (3) our fine-grained degradation representation via image reconstruction. The OneRestore extractor, optimized for classification, is insensitive to degradation strength or location, leading to poorer results. Our method surpasses the non-blind ”Classes” baseline by 0.11 dB PSNR.
Fig. 8 visualizes the t-SNE embeddings of the extracted degradation features. Significant gaps exist between different degradation types, indicating clear separability. Notably, although JPEG and noise share the same content during synthesis, the t-SNE visualization reveals that their separability is primarily driven by degradation information rather than image content.
@@
TABLE V: Ablation Study on Different Degradation Extraction and Utilization Methods.
@@
Comparison Projects Method PSNR SSIM Degradation Extraction Classes 27.58 0.892 OneRestore* 27.56 0.891 Ours 27.69 0.893 Degradation Insertion Concat 26.96 0.882 Attention 27.39 0.882 Ours 27.69 0.893
Table VI presents the ablation study on global information supplementation. We compare our Prompt-C with Mamba’s bi-directional scanning scheme. Compared to Mamba’s bi-directional scan, our Prompt-C not only achieves slightly better performance but is also more efficient, reducing GFLOPs by nearly 37% (Table VI). This highlights its effectiveness as a lightweight global information supplementation method.
@@
TABLE VI: Ablation Study on Different Methods for Global Information Supplementation.
@@
Method PSNR SSIM Params(M) GFLOPs Bi-direction Scan 27.66 0.893 31.43 244.3 Prompt Matrix C (Ours) 27.69 0.893 29.35 153.3

### v3

@@ III-D Degradation-aware Prompt State Space Model
To address the limitations of SSM-based image restoration, specifically the insufficient use of global context and the lack of effective degradation-aware mechanisms, we propose a novel Degradation-aware Prompt State Space Model (DP-SSM) to effectively adapt SSM to All-in-One image restoration. Specifically, we propose a novel Degradation-aware Prompt Selective Scan, which incorporates degradation information into the State Space Model. This is achieved by leveraging a Degradation Embedding extracted from a pre-trained Degradation Extractor to provide supplementary degradation-related contextual information that the traditional SSM cannot capture. In addition, we modulate the parameter matrices in the state-space equations to establish a degradation-aware mechanism.
This mechanism enables SSM to adapt to multi-task learning and handle task competition. To better recover local high-frequency details, we design a lightweight High-frequency Enhancement Block (HEB). This component ensures the preservation of high-frequency details and enhances the overall model performance.
@@
where $init()$ is an initialization method that does not depend on the input $x$. However, this input-dependent design is agnostic to the degradation type. It processes high-frequency noise and low-frequency blur with the same initialization prior, lacking the ability to structurally adapt the system dynamics for specific restoration tasks. This motivates the introduction of a degradation-aware prompt mechanism, which guides the state-space modeling by incorporating specific degradation cues.
@@
$\overline{B}_{dp}$ $=(\Delta_{dp}A)^{-1}(\exp(\Delta_{dp}A)-I)\cdot\Delta_{dp}B_{dp},$
@@
The proposed modulation mechanism fundamentally alters the underlying system dynamics, effectively recalibrating the trade-off between memory inertia and input sensitivity. Central to this adaptation is the step size $\Delta_{dp}$, whose learned distribution shown in Fig. 5 reveals a distinct physical correspondence to the degradation characteristics. For high-frequency degradations such as noise and JPEG artifacts, the model autonomously converges to a minimal $\Delta_{dp}$ regime. As $\Delta_{dp}$ approaches zero, the discretized transition matrix $\overline{{A}}_{dp}$ approximates the identity matrix while the input gain $\overline{{B}}_{dp}$ diminishes. This configuration transforms the SSM into a stable low-pass filter, creating a high-inertia state that effectively suppresses instantaneous pixel variance by prioritizing historical context over the unreliable current observation. In contrast, inputs characterized by low-frequency or structural deficiencies, such as blur and low-light conditions, trigger a marked increase in the learned $\Delta_{dp}$. This shift transitions the system into a high-gain regime, amplifying the magnitude of $\overline{{B}}_{dp}$ to enhance sensitivity towards weak input signals and facilitate the recovery of attenuated local structures. Complementing this dynamic discretization, the concurrent modulation of B{B} acts as a gating mechanism for input observations, while the modulation of C{C} supplements the feature readout with global degradation cues.
Figure 5: Distribution of the learned parameter $\Delta_{dp}$ across different degradation types in the first 20% of DPSS layers. The model adaptively assigns smaller step sizes for noise/jpeg and larger step sizes for blur/low-light, confirming physically interpretable dynamic modulation.
Through this strategy, the transformation matrix A{A} and input matrix B{B} are endowed with degradation-aware capabilities, allowing them to adaptively regulate memory inertia and input sensitivity according to the specific degradation severity. However, the standard SSM remains constrained by its strictly causal nature, where the $i$-th pixel can only access information from the preceding i1i-1 pixels, resulting in a restricted global receptive field. Although existing methods attempt to alleviate this via parallel multi-directional scanning, such approaches incur high computational costs and yield limited performance gains in low-level vision tasks [9]. By retaining the efficient single-directional scanning and instead modulating matrix C{C} to inject global degradation information, our method achieves effective global degradation awareness with significantly lower computational overhead.
@@
High-frequency Enhancement Block. All-in-One image restoration necessitates the recovery of image components distributed across diverse frequency bands. While our proposed Degradation-aware Prompt State Space Model enables dynamic adaptation to degradation characteristics and partially mitigates inter-task conflicts, the joint optimization of heterogeneous degradations inevitably induces an implicit bias toward low-frequency structures. As illustrated in Fig. 6, this spectral bias results in suboptimal performance regarding high-frequency details, such as edges and textures, compared to degradation-specific models. To address this limitation, we introduce a lightweight High-Frequency Enhancement Block (HEB) as a complementary module to explicitly facilitate the restoration of local textures.
@@ III-E Loss Function
where $\hat{O}$ and $O$ denote the model output and the ground truth, respectively. The parameters $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are balancing factors, which we set to 1, 0.5 and 0.001, respectively.
@@
TABLE I: Quantitative Comparison on Seven Datasets for All-in-One Image Restoration. The best and second-best results are shown in bold and underline, respectively. (Format: PSNR $\uparrow$  / SSIM $\uparrow$  / LPIPS $\downarrow$).
Method Lowlight Snow Rain Haze Jpeg Noise Blur Average MPRNet 19.71 / .659 / .122 30.94 / .938 / .029 26.27 / .912 / .068 25.63 / .947 / .023 30.51 / .938 / .024 28.03 / .897 / .057 26.07 / .869 / .077 26.74 / .880 / .057 MIRNet 19.81 / .668 / .123 31.04 / .941 / .028 25.77 / .907 / .073 22.26 / .882 / .058 30.54 / .939 / .024 28.14 / .900 / .057 26.32 / .873 / .070 26.27 / .873 / .062 NAFNet 19.49 / .650 / .117 30.33 / .935 / .030 26.04 / .906 / .068 25.63 / .946 / .023 30.20 / .936 / .025 27.80 / .895 / .051 26.13 / .870 / .075 26.52 / .877 / .056 Restormer 19.99 / .670 / .113 31.67 / .944 / .025 26.91 / .919 / .061 26.18 / .953 / .020 30.56 / .939 / .024 28.17 / .890 / .056 26.62 / .880 / .063 27.16 / .886 / .052 MambaIR 19.81 / .664 / .115 30.43 / .935 / .032 25.52 / .901 / .078 25.90 / .944 / .024 30.38 / .937 / .025 27.91 / .895 / .058 26.07 / .869 / .077 26.57 / .878 / .058 PromptIR 19.90 / .669 / .116 31.52 / .943 / .024 26.80 / .918 / .061 26.13 / .945 / .022 30.57 / .939 / .024 28.18 / .901 / .054 26.36 / .874 / .072 27.07 / .884 / .053 IDR 19.68 / .663 / .115 31.16 / .941 / .027 26.39 / .914 / .065 26.11 / .948 / .023 30.51 / .939 / .024 28.10 / .899 / .054 26.29 / .873 / .073 26.89 / .883 / .054 NDR-Restore 19.79 / .635 / .206 30.00 / .902 / .052 26.00 / .813 / .158 27.00 / .943 / .025 30.19 / .872 / .056 27.78 / .786 / .133 25.77 / .791 / .193 26.64 / .820 / .117 OneRestore 19.77 / .657 / .127 29.95 / .931 / .034 25.37 / .900 / .080 27.74 / .962 / .016 30.10 / .934 / .026 27.65 / .890 / .060 25.90 / .866 / .082 26.64 / .877 / .061 MoCEIR 19.92 / .674 / .109 31.22 / .943 / .025 26.56 / .920 / .051 28.06 / .963 / .014 30.51 / .939 / .023 28.13 / .900 / .053 26.34 / .874 / .069 27.25 / .888 / .049 AdaIR 19.99 / .672 / .102 31.94 / .923 / .023 27.01 / .922 / .056 26.61 / .952 / .019 30.53 / .939 / .024 28.19 / .900 / .055 26.61 / .879 / .065 27.26 / .887 / .049 DPMambaIR 20.04 / .680 / .109 32.36 / .949 / .022 27.07 / .926 / .055 28.13 / .963 / .016 30.57 / .939 / .024 28.25 / .902 / .051 27.43 / .894 / .048 27.69 / .893 / .047 DPMambaIR-L 20.06 / .677 / .108 32.31 / .950 / .022 27.11 / .925 / .056 28.39 / .963 / .016 30.55 / .939 / .024 28.21 / .902 / .051 27.45 / .895 / .045 27.73 / .893 / .046
@@ IV-A Experimental Settings
Training Details. To ensure effective degradation-aware modeling, we adopt a two-stage training strategy. In the first stage, we pre-train the Degradation Extractor using the same mixed-degradation dataset employed for the All-in-One task. The training pipeline consists of the Degradation Extractor, a Content Extractor, and a Degradation Reconstructor. Specifically, we utilize the pre-trained Restormer from DegAE [25] as the Content Extractor with its parameters frozen, while optimizing only the Degradation Extractor and Reconstructor. This pre-training phase is conducted for 100,000 iterations with a batch size of 1, using input patches cropped to $256\times 256$. We employ the AdamW [20] optimizer ($\beta_{1}=0.9$, $\beta_{2}=0.999$, weight decay $1e^{-4}$) with an initial learning rate of $1e^{-4}$, which is gradually reduced to $1e^{-6}$ through cosine annealing.
In the second stage, the pre-trained Degradation Extractor is integrated into the DPMambaIR framework with its parameters frozen to provide stable guidance. Building on previous works [57, 39, 9], we train the main restoration network using the same AdamW configuration for 300,000 iterations. The initial learning rate is set to $3e^{-4}$ and decayed to $1e^{-6}$ via cosine annealing. Following [57], we employ a progressive training strategy, cropping image patches of size $192\times 192$. Horizontal and vertical flips are applied for data augmentation. Our experiments are conducted on a single NVIDIA RTX A100 GPU and implemented using the PyTorch platform.
@@
Evaluation Metrics. Following previous works [57, 9], we employ Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index Measure (SSIM) to quantitatively evaluate restoration quality on RGB channels. Additionally, we utilize the Learned Perceptual Image Patch Similarity (LPIPS) to assess the perceptual consistency of the restored images.
@@ IV-B Quantitative Results
All-in-one Image Restoration. We evaluated two versions of our method for comparison: the normal version (DPMambaIR) and a large version (DPMambaIR-L). The normal version adopts the degradation extractor illustrated in Fig. 3(c), while the large version uses the architecture shown in Fig. 3(b). For subsequent evaluations, we primarily adopt the normal version unless otherwise specified. We compare our method against eleven state-of-the-art image restoration approaches. As shown in Table I, for the All-in-One image restoration task, our method achieves the best PSNR and SSIM on the aforementioned mixed datasets, outperforming existing methods. Specifically, our method achieves a PSNR of 27.69 dB and an SSIM of 0.893, surpassing AdaIR by 0.43 dB and 0.011, respectively. Additionally, DPMambaIR-L achieves a PSNR of 27.73 dB and an SSIM of 0.893, slightly outperforming the normal version.
@@ IV-C Qualitative Results
IV-D Model Complexity and Efficiency Analysis
To accurately evaluate the practical applicability of DPMambaIR, we conducted a comprehensive efficiency comparison against state-of-the-art All-in-One methods. We measured the number of parameters, peak memory consumption, GFLOPs, and average inference time. All evaluations were performed on a single NVIDIA RTX A100 GPU with an input tensor size of $1\times 3\times 256\times 256$. The statistics for DPMambaIR include the overhead of the Degradation Extractor.
As summarized in Table IV, DPMambaIR achieves the best performance while maintaining comparable computational overhead. Compared to the second-best method, AdaIR, our model delivers a significant performance gain of 0.43 dB with similar parameters and GFLOPs.
TABLE IV: Comparison of complexity and efficiency against SOTA methods. Metrics: Parameters (M), Peak Memory (MB), GFLOPs, and Inference Time (ms) on $256\times 256$ inputs.
Method Params (M) Mem. (MB) FLOPs (G) Inf Time (ms) PSNR (dB) Restormer 26.13 676.24 154.88 84.283 27.16 IDR 36.28 1433.24 270.76 81.977 26.89 PromptIR 35.59 720.10 172.71 90.964 27.07 MoCEIR 25.35 404.48 97.74 93.232 27.25 AdaIR 28.78 686.40 161.76 102.903 27.26 DPMambaIR 29.35 718.76 153.33 90.869 27.69
IV-E Ablation Study
We conduct ablation studies to evaluate our key designs: the degradation-aware prompts and High-frequency Enhancement Block (HEB) (Table V), degradation extraction and utilization methods (Table VI), degradation embedding dimension (Table VII) and the global information supplementation strategy (Table VIII).
Table V demonstrates the effectiveness of our proposed core modules. First, the High-frequency Enhancement Block (HEB) alone improves the baseline by 0.29 dB (Model (a)), validating its efficacy in enhancing local details. Second, to verify the importance of the degradation-aware mechanism, we conduct a detailed breakdown of the prompt components. Notably, employing $P_{\Delta}$ alone (Model (c)) achieves 27.58 dB, which slightly outperforms the combination of $P_{B}$ and $P_{C}$ (Model (b), 27.56 dB). This observation underscores that dynamically modulating the discretization step $\Delta$, which fundamentally alters the state transition dynamics, is the most critical factor for adapting to diverse degradation types. Furthermore, combining $P_{\Delta}$ with $P_{B}$ or $P_{C}$ yields consistent improvements, and integrating all prompt modules (Model (f)) pushes the PSNR to 27.67 dB. Finally, the full model (Model (g)) demonstrates the synergy between prompt-based modulation and frequency enhancement, achieving the best performance of 27.69 dB.
@@ IV-E Ablation Study
Table VI presents the results for degradation extraction and utilization methods. For degradation extraction, we compare three approaches: (1) “Classes”, which provides manually specified degradation types and can be regarded as a 100% accurate non-blind setting; (2) a pre-trained degradation classifier from OneRestore [11]; and (3) our fine-grained degradation representation via image reconstruction. The OneRestore extractor, optimized for classification, is insensitive to degradation strength or location, leading to poorer results. Our method surpasses the non-blind ”Classes” baseline by 0.11 dB PSNR. Fig. 9 visualizes the t-SNE embeddings of the extracted degradation features. Significant gaps exist between different degradation types, indicating clear separability. Notably, although JPEG and noise share the same content during synthesis, the t-SNE visualization reveals that their separability is primarily driven by degradation information rather than image content.
@@
For the training strategy, compared with jointly fine-tuning the degradation extractor together with the backbone network, freezing the parameters of the degradation extractor provides a more stable supply of degradation information. This design choice leads to more consistent guidance during restoration and results in a clear performance gain of 0.26 dB in PSNR, demonstrating the advantage of decoupling degradation representation learning from the restoration optimization process.
TABLE V: Extended ablation studies on the proposed core modules. We investigate the individual and combined contributions of the prompt components ($P_{\Delta},P_{B},P_{C}$) and the High-frequency Enhancement Block (HEB).
Method $P_{\Delta}$ $P_{B}$ $P_{C}$ HEB PSNR SSIM Baseline 26.96 0.884 (a) 27.25 0.887 (b) 27.56 0.891 (c) 27.58 0.892 (d) 27.60 0.892 (e) 27.61 0.892 (f) 27.67 0.893 (g) 27.69 0.893
@@
TABLE VI: Ablation Study on Different Degradation Extraction and Utilization Methods.
@@
Module Method PSNR SSIM Degradation Extraction Classes 27.58 0.892 OneRestore* 27.56 0.891 Ours 27.69 0.893 Degradation Insertion Concat 26.96 0.882 Attention 27.39 0.882 Modulation (Ours) 27.69 0.893 Training Strategy Fine-tuning 27.43 0.888 Frozen (Ours) 27.69 0.893
TABLE VII: Ablation study on the dimension of degradation embedding ($D_{emb}$). We evaluate model complexity (Params, GFLOPs), auxiliary task performance (Rec. PSNR, Cls. Acc.), and final All-in-One restoration quality (AIO PSNR). The default configuration is gray.
$D_{emb}$ Params FLOPs Rec. PSNR Cls. Acc. AIO. PSNR (M) (G) (dB) (%) (dB) 256 28.71 153.33 21.51 72.0 27.55 384 29.03 153.33 24.53 84.3 27.63 512 29.35 153.33 25.67 92.0 27.69 1024 30.62 153.33 25.78 92.6 27.69
@@
TABLE VIII: Ablation Study on Different Methods for Global Information Supplementation.
@@
Method PSNR SSIM Params (M) GFLOPs Bi-direction Scan 27.66 0.893 31.43 244.3 Prompt-C (Ours) 27.69 0.893 29.35 153.3
Table VII presents the ablation study on the Degradation Embedding dimension ($D_{emb}$). Given that the Degradation Extractor serves as the sole source of degradation priors, the completeness and distinctiveness of the extracted information are paramount. To quantitatively evaluate the quality of the learned embeddings, we introduce two direct indicators: (1) Degradation Reconstruction PSNR (Rec. PSNR), calculated between the reconstructed degraded image and the target degraded input, which measures information integrity; and (2) Classification Accuracy (Cls. Acc.), which assesses the semantic separability of the degradation types. For the classification metric, we appended a simple MLP head to the frozen degradation extractor and trained it for only 1,000 iterations to map the embeddings to the seven degradation classes. As observed in Table VII, increasing $D_{emb}$ from 256 to 512 yields substantial gains: Rec. PSNR improves from 21.51 dB to 25.67 dB, and Cls. Acc. surges from 72.0% to 92.0%. This indicates that higher-dimensional embeddings effectively encode both the pixel-level details and the semantic categories of degradations. These improvements directly translate to the final restoration task, boosting AIO. PSNR to 27.69 dB. However, expanding to 1024 yields marginal returns despite higher costs. Thus, 512 is selected as the optimal balance.
Table VIII presents the ablation study on global information supplementation. We compare our Prompt-C with Mamba’s bi-directional scanning scheme. Compared to Mamba’s bi-directional scan, our Prompt-C not only achieves slightly better performance but is also more efficient, reducing GFLOPs by nearly 37%. This highlights its effectiveness as a lightweight global information supplementation method.
IV-F Generalization Capability and Failure Boundary Analysis
To investigate the robustness of DPMambaIR beyond the training distribution, we conducted Out-of-Distribution (OOD) evaluations on two unseen degradation types: Raindrop and Pixelation. Importantly, these evaluations were performed by directly applying the pre-trained model to the test images without any additional training or fine-tuning. This rigorous setting serves to delineate the generalization boundaries of our Degradation-Aware Prompt mechanism.
Robustness on Unseen Degradations. We evaluated the model on the Raindrop dataset[40] (58 images) and a synthetically generated Pixelated BSD68 dataset[29] (block size 2). As summarized in Table IX, DPMambaIR exhibits superior generalization, securing top-tier performance across all metrics. Specifically, on the Pixelation task, it achieves a PSNR of 23.35 dB, surpassing the second-best AdaIR by a significant margin of 1.96 dB. We attribute this robustness to our regression-based embedding strategy. Unlike static prompt approaches that risk model collapse on undefined inputs, DPMambaIR projects unseen degradations into a continuous latent space. This mechanism enables the model to approximate unknown corruption patterns within a learned manifold, preventing inference failure and yielding visually plausible results. A notable phenomenon is further observed in a specific low-light sample from the BSD68 dataset. When processing the pixelated version of this image, DPMambaIR automatically performs brightness enhancement alongside depixelation, capturing and correcting the underlying low-light degradation. In contrast, competing methods fail to address this compound issue. This capability underscores the potential of DPMambaIR in handling mixed degradations, suggesting that future work can further explore real-world restoration via more complex, fine-grained coupled degradation modeling.
Failure Boundary Analysis. Despite these promising results, DPMambaIR is not without limitations. We observed a distinct performance bottleneck when processing images with heavy raindrops. As visually analyzed in Fig. 10, while the model effectively removes light raindrops, it struggles with large, opaque water droplets that completely occlude the background. We attribute this to the physical nature of the degradation: removing heavy, opaque occlusions requires generative hallucination (inpainting) to synthesize missing content, whereas DPMambaIR operates on restoration cues. Future work could explore incorporating generative priors, such as Diffusion Models, to address such severe occlusions.
TABLE IX: Quantitative comparison on unseen degradation types (OOD Testing). The best and second-best results are highlighted in bold and underline.
Method Pixelation Raindrop
(PSNR / SSIM / LPIPS) (PSNR / SSIM / LPIPS)
MambaIR 20.80 / 0.7601 / 0.0568 19.29 / 0.7992 / 0.1051
PromptIR 19.69 / 0.7479 / 0.0582 19.26 / 0.7775 / 0.1146
IDR 20.64 / 0.7567 / 0.0494 17.40 / 0.7609 / 0.1276
OneRestore 19.75 / 0.7415 / 0.0628 19.75 / 0.7792 / 0.1100
AdaIR 21.39 / 0.7676 / 0.0512 19.07 / 0.7722 / 0.1225
MoCEIR 19.89 / 0.7355 / 0.0544 21.46 / 0.8254 / 0.0919
DPMambaIR 23.35 / 0.8003 / 0.0551 21.63 / 0.8337 / 0.0932
Figure 10: Visual analysis of generalization and failure boundaries. (a) Success Cases: The top row demonstrates restoration on unseen Pixelation, where DPMambaIR concurrently removes artifacts and enhances inherent low-light details. The middle row shows effective removal of Light Raindrop. (b) Failure Case: The bottom row illustrates the boundary on Heavy Raindrop, where the model fails to hallucinate the background behind opaque occlusions.